OpenStudy (anonymous):

Solve this linear equation: 4c+23=9c-7 Please show how to solve step by step.

OpenStudy (anonymous):

4c - 9c = -7 - 23 -5c = -30 c = 6

OpenStudy (anonymous):

4c-9c=-7-23, -5c=-30, c=6

OpenStudy (anonymous):

You can check the result by plugging in 6 back into the equation: 4(6) + 23 = 9(6) - 7 47 = 47 so your solutions is true.